OpenStudy (anonymous):

How many turns does a wheel need with radius of 14 inches to across a 1 mile track?

OpenStudy (anonymous):



OpenStudy (anonymous):

circumference of wheel = \[2 \pi r\] r = 14 inches So the number of turns to cover one mile (63360 inches) will be \[\left( 63360 \right) \div \left( 2 \pi \times 14 \right)\] = 360 (approx)

OpenStudy (anonymous):

2640

OpenStudy (anonymous):

Sorry there has been a calculation mistake, the answer comes out to be 720.28 but still it isnt 2640.